# cosmICweb: Cosmological Initial Conditions for Zoom-in Simulations in the Cloud

### v1

@@ III.1 Minimal bounding ellipsoids in cosmICweb
We outline the computed minimum bounding ellipsoids for the selected halos in Figure 2. Note that the illustration is only a 2-dimensional slice through the center of the ellipsoid, hence the ellipsoids appear to be a bad fit in some cases since the 3-dimensional structure is not visible. Nevertheless, we clearly notice that halos in isolated environments tend to have more “efficient” ellipsoids, with a higher fraction of particles within the proto-halo collapsing into the traceback-radius. We discuss the ellipsoid efficiency in a more statistical analysis in Section V.1.
@@
For the simulations provided by the cosmICweb team (labeled “primary” in Table 1), we provide minimal bounding ellipsoids for free halos (i.e., no subhalos) at every snapshot and for four different traceback-radii $R_{\mathrm{tb}}=1,2,4$, and 10 $R_{\mathrm{vir}}$ if the number of particles within $R_{\mathrm{tb}}$ is larger than the threshold $N^{e}_{\min}$. The EAGLE simulations provide minimum bounding ellipsoids at $R_{\mathrm{tb}}=1,2,4$, and 8 $R_{200c}$ for all FoF groups at each snapshot. The ellipsoids are computed using Khachiyan’s algorithm (Khachiyan, 1996; Kumar & Yildirim, 2005), which iteratively optimizes the volume of the ellipsoid under the constraint that all particles are contained. We account for periodic boundary conditions by moving all particles to the same local coordinate system centered at the proto-halo center of mass. This requires that no axis of the proto-halo patch is longer than half the box size, which generally should not occur and for which a zoom-simulation would provide only marginal benefits.
IV Database architecture and web application
@@ IV.1 Architecture
Finally, the cosmicweb-music command-line interface (CLI) is a convenient tool that allows users to download initial condition configuration files directly via the command line. By using the CLI, users can avoid the cumbersome process of downloading files locally and then re-uploading them to a headless compute cluster can be avoided.
For the interaction between these four components to function seamlessly, requests to and responses from the API servers must adhere to a specified protocol, which we further discuss in Appendix A. More detailed and up-to-date documentation is also available directly on the cosmICweb website [^5] https://cosmicweb.eu/documentation/api. How these requirements are best implemented on the API server side needs to be evaluated on a case-by-case basis, considering factors such as the available data, its storage, and the server infrastructure. We discuss the implementation of the API server used for the “primary” simulations in more detail at the end of this section.
IV.2 Web Interface
@@
Table 2: List of currently available selectors, grouped by halo properties, galaxy properties, and further parameters including properties of available proto-halo patches. [L] and [E] mark the features that are currently supported in the locally hosted “primary” simulations and the Eagle simulations.**footnotetext: the primary simulations currently do not support $M_{500b}$, $M_{2500b}$ and the corresponding radii and concentration parameter.footnotetext: the Eagle simulations currently support $M_{\mathrm{200b}}$, $M_{\mathrm{200c}}$, $M_{\mathrm{500c}}$, $M_{\mathrm{2500c}}$, and $R_{\mathrm{200c}}$.
@@
halo $M_{200c}$ / $M_{500c}$ / $M_{2500c}$ $h^{-1}\mathrm{M}_{\odot}$ Mass within $\rho>=\Delta\rho_{c}$ contrast
@@
$\lambda_{\mathrm{Peebles}}$ Spin parameter (Peebles definition, Peebles (1971))
$\lambda_{\mathrm{Bullock}}$ Spin parameter (Bullock definition, Bullock et al. (2001))
@@
[…]
last major merger redshift of last major merger
@@
$\epsilon$ Efficiency of the Lagrangian minimum bounding ellipsoid $\epsilon=n_{\mathrm{part}}/V_{\mathrm{ellipsoid}}$
include substructure boolean option to include substructure in the search
Once a query is submitted, the API server returns a list of matching halos. The properties of these halos can be displayed in a table and visualized on two-dimensional scatter plot. From the results panel, users can open halos in the Halo Explorer for a more detailed overview, group and store them as collections, and directly download as initial condition files for resimulation.
@@
IV.3 API implementation for the cosmICweb datasets
@@
V Discussion and Validation
In this section, we provide a quantitative discussion of the minimum bounding ellipsoids available on cosmICweb and demonstrate how re-simulation targets based on “well-behaved” ellipsoid potentially leads to a biased selection of halos. Additionally, we present two zoom simulations of halos in our catalogs as a proof-of-concept to demonstrate the capability of cosmICweb.
V.1 Efficiency distribution of minimum bounding ellipsoids
Figure 7: Relative change of mass of Lagrangian proto-halo patches (top) and the efficiency parameter (bottom) with increasing $R_{\mathrm{tb}}$ for halos in the mass bins detailed in Table 3. For the mass increase, we separately plot the change in $M_{\mathrm{ellipsoid}}$ (dashed) and $M_{\mathrm{particles}}$ (solid) separately. In addition, we subdivide two mass bins according to the environmental parameter $D_{1,1}$, shown on the right.
mass range # halos simulation # particles
$M_{\mathrm{vir}}$ $h^{-1}\mathrm{M}_{\odot}$ ] per halo
$(3,6)\times 10^{10}$ 140736 150MPC 110 – 220
$(1,2)\times 10^{11}$ 51407 150MPC 370 – 740
$(1,2)\times 10^{12}$ 52030 300MPC 470 – 930
$(1,2)\times 10^{13}$ 6449 300MPC 4670 – 9340
$(1,2)\times 10^{14}$ 17803 Rhapsody_
PlanckCosmo
1250 – 2500
Choosing a larger traceback-radius increases the volume of the minimal bounding ellipsoid. The exact growth of the Lagrangian ellipsoid depends on the mass and the environment of the halo. To analyze the mass growth of the proto-halo patches at different traceback radii, we compare halos from five mass bins, detailed in Table 3. Additionally, we split the $10^{11}$ $h^{-1}\mathrm{M}_{\odot}$ and $10^{13}$ $h^{-1}\mathrm{M}_{\odot}$ bins into three different environmental regimes, determined by the distance to the closest more massive halo $D_{1,1}$ (see Table 2). The results are shown in Figure 7.
V.2 Biased target selection by efficiency constraints
In order to measure these effects, we select halos with Lagrangian efficiencies below the 10th and above the 90th percentile from the $M_{\mathrm{vir}}\in[2,4]\times 10^{12}$ $h^{-1}\mathrm{M}_{\odot}$ mass bin of the 300MPC simulation at $z=0$. We then compute the distributions of a selection of external and intrinsic parameters and compare it to the distribution of all halos in that mass bin. The chosen parameters include the normalized distance to the closest more massive halo $D_{1,1}$, the relative mass increase between $1$ and $10R_{\mathrm{vir}}$ ($M_{10,1}$), the concentration parameter $c_{\mathrm{vir}}$, the spin parameter $\lambda_{\mathrm{Peebles}}$ (cf. Table 2), and the ellipticity of the minimum bounding ellipsoid $e=(a_{3}-a_{1})(2\sum a_{i})$, where $a_{i}=\lambda_{i}^{-1/2}$ are the semi-axes of the ellipsoid related to the eigenvalues $\lambda_{1}\geq\lambda_{2}\geq\lambda_{3}$ of the shape matrix (cf. Equation 2). The first two parameters measure the local clustering, while the following two are intrinsic properties of the halo. The ellipticity measures how spherical (and therefore well-behaved) the proto-halo patch is. Merger events and strong tidal fields are expected to distort the proto-halo patch, leading to a more elliptical and less efficient bounding ellipsoid.
V.3 Example Simulations
As a proof-of-concept, we present two zoom simulations originating from the Eagle RefL0100N1504 volume. In addition to demonstrating the capability of cosmICweb, these examples also serve as a check of the Panphasia white-noise field integration into music. One of the challenges to incorporate Panphasia is that (currently) music requires power-of-two box resolutions due to the implementation of a multigrid Poisson solver to compute LPT terms (see Hahn & Abel, 2011). This constraint requires a Fourier-space remapping of Panphasia fields. Specifically, we generate the Panphasia fields at the native resolution, e.g., $3\times 2^{\ell}$ where $\ell\in\mathbb{N}$ is some subdivision level, and then cut out the modes representable at the next smaller power of two (in the example $2^{\ell+1}$). This naturally introduces some small discrepancies and warrants an investigation how well we can reproduce existing zoom simulations with music and its Panphasia integration.
@@ V.3 Example Simulations
Note that music allows to specify a minimum refinement level for the entire simulation, levelmin, which sets the number and mass of the coarsest particles in the zoom simulation, and a minimum refinement level when generating the initial conditions levelmin_TF. The particle displacements and velocities are later averaged down to levelmin. We present simulations varying both. This tests the dependence on the coupling of small-scale modes into the high-res regions, i.e. the accuracy of the zoom approach itself as well as its implementation in music, and the impact of a coarsely represented tidal field on the non-linear evolution during the simulation.
Milky-Way sized halo
@@
Figure 9: Visual comparison of the Au6 halo (Grand et al., 2017) between the reference simulation and three zoom simulations (cf. Figure 10). We calculate the surface density using a Delaunay tessellation field estimator (Rangel et al., 2016). The white circle indicates $R_{\mathrm{200c}}$ determined in the reference simulation.
Figure 10: Density profile comparison for the Au6 halo (Grand et al., 2017). We compare three zoom simulations initialized with music using proto-halo patches from cosmICweb against the reference simulation by Grand et al. (2017) that has been initialized with Panphasia. The proto-halo patches have been calculated at a traceback radius of $4R_{\mathrm{200c}}$. In addition to the total matter density profiles (solid) we highlight the contribution of low resolution particles outside the refinement region separately (dashed). The lower panel shows the relative difference of the total matter density profile with respect to the reference solution.
@@
simulation (7,8,10) (7,10,10) (7,10,12) Ref.
$m_{p}$ $10^{5}\mathrm{M}_{\odot}$] $364$ $364$ $5.69$ $2.34$
$\epsilon_{\rm comoving}$ $\mathrm{kpc}$] 3.26 3.26 0.813 0.739
$M_{200,c}$ $10^{10}\mathrm{M}_{\odot}$] 105.3 108.1 105.2 105.4
$M_{500,c}$ $10^{10}\mathrm{M}_{\odot}$] 83.2 83.3 83.0 82.0
$M_{2500,c}$ $10^{10}\mathrm{M}_{\odot}$] 48.0 48.0 49.8 48.5
@@
The first example is the Au6 halo from the Auriga Project (Grand et al., 2017, 2024). This Milky Way mass halo has previously been studied to analyze formation mechanisms of disc galaxies. We perform gravity only zoom simulations at different levels of refinement of the Au6 Lagrangian patch, calculated at a traceback radius $R_{\mathrm{tb}}=4R_{\mathrm{200c}}$ at redshift 0. The halo can be found in the publication section on cosmICweb  [^7] Auriga 6: https://cosmicweb.eu/publication/auriga-halos.
We re-simulated this halo at two different mass resolutions (levelmax=10 and 12). In addition, we once generated the initial noise field at the full resolution followed by averaging in the coarse regions (levelmin_TF=levelmax=10 : (7,10,10)), and using the zoom technique (levelmin_TF << levelmax: (7,8,10) and (7,10,12)). The re-simulation was carried out using the swift $N$-body code (Schaller et al., 2024). The respective mass resolution of all runs along with the force softening are indicated in Table 4. We compare these halos with the redshift zero output of a reference calculation, which is an original Auriga level 4 gravity only simulation run with gadget. The initial conditions for the reference were created by the ic_gen code (Jenkins, 2010, 2013) using a $1536^{3}$ cell patch from the Panphasia field. The nearest smaller representable resolution in music is levelmin_TF=10 with $2^{10}=1024$. We determined halo properties using the rockstar halo finder (Behroozi et al., 2013a) and list the recovered main properties of the halos in Table 4, as well as show the projected mass distribution in Figure 9 and spherically averaged differential mass density profiles in Figure 10.
@@
Regarding global properties, we find that the halo mass $M_{\mathrm{200,c}}$ is recovered at the level of better than 1 per cent for two re-simulations, while the third, $(7,10,10)$, is somewhat more discrepant at $\sim 3$ per cent. The virial radius always agrees to better than one per cent. Properties of the inner mass distribution (concentration $c_{200c}$ and mass $M_{2500c}$) of the halos are slightly more discrepant but still consistent at the few per cent level. Note that the rockstar measurement of the mass and other properties of the reference run is slightly discrepant from the values reported in Grand et al. (2024) using subfind.
@@
The visual impression from Figure 9 is similar in the overall shape of the halo, also positions of some halos outside the virial radius are in excellent agreement. Interestingly, the most massive substructure is however at different locations in all of the runs — likely a consequence of it being very sensitive to any perturbations of the initial conditions.
@@
Finally, the density profiles shown in Figure 10 show agreement at the few per cent level at all radii up to the radius affected by finite force and mass resolution and out to the maximum radius of the zoom region of $4R_{200c}$.
Group-sized halo

### v3

@@ 3.1 Minimal bounding ellipsoids in cosmICweb
We outline the computed minimum bounding ellipsoids for the selected halos in Figure 2. Note that the illustration is only a 2-dimensional slice through the center of the ellipsoid, hence the ellipsoids appear to be a bad fit in some cases since the 3-dimensional structure is not visible. Nevertheless, we clearly notice that halos in isolated environments tend to have more “efficient” ellipsoids, with a higher fraction of particles within the proto-halo collapsing into the traceback-radius. We discuss the ellipsoid efficiency in a more statistical analysis in Section 3.2. Also note that some of the Lagrangian patches at low traceback radii contain holes, in particular for halos in more populated regions (bottom panels). These holes originate from particles that have previously traversed the halo and are now outside the overdensity radius. Sampling those holes with coarse particles would could lead to artifacts as they traverse the halo; however, using a minimum bounding ellipsoids as the refinement region will ensure that those particles that pass through the halo are also sampled in high resolution.
@@
For the simulations provided by the cosmICweb team (labeled “primary” in Table 1), we provide minimal bounding ellipsoids for free halos (i.e., no subhalos) at every snapshot and for four different traceback-radii $R_{\mathrm{tb}}=1,2,4$, and 10 $R_{\mathrm{vir}}$ if the number of particles within $R_{\mathrm{tb}}$ is larger than the threshold $N^{e}_{\min}$. The EAGLE simulations provide minimum bounding ellipsoids at $R_{\mathrm{tb}}=1,2,4$, and 8 $R_{200,c}$ for all FoF groups at each snapshot. The ellipsoids are computed using Khachiyan’s algorithm (Khachiyan, 1996; Kumar & Yildirim, 2005), which iteratively optimizes the volume of the ellipsoid under the constraint that all particles are contained. We account for periodic boundary conditions by moving all particles to the same local coordinate system centered at the proto-halo center of mass. This requires that no axis of the proto-halo patch is longer than half the box size, which generally should not occur and for which a zoom-simulation would provide only marginal benefits.
Note that ellipsoids derived from small traceback radii, in particular from lower mass halos that are not well-sampled by particles, can cause contamination of the zoom region by coarse-resolution particles, leading to potential numerical artifacts. We recommend to choose ellipsoids from larger trace-back radii or to run an intermediate zoom simulation to refine the zoom region in this case.
3.2 Efficiency distribution of minimum bounding ellipsoids
Figure 4: Relative change of mass of Lagrangian proto-halo patches (top) and the efficiency parameter (bottom) with increasing $R_{\mathrm{tb}}$ for halos in the mass bins detailed in Table 2. For the mass increase, we plot the change in $M_{\mathrm{ellipsoid}}$ (dashed) and $M_{\mathrm{particles}}$ (solid) separately. In addition, we subdivide two mass bins according to the environmental parameter $D_{1,1}$, shown on the right.
mass range # halos simulation # particles
$M_{\mathrm{vir}}$ [$h^{-1}\mathrm{M}_{\odot}$ ] per halo
$(1,2)\times 10^{11}$ 51407 150MPC 370 – 740
$(1,2)\times 10^{12}$ 52030 300MPC 470 – 930
$(1,2)\times 10^{13}$ 6449 300MPC 4670 – 9340
$(1,2)\times 10^{14}$ 17803 Rhapsody_
PlanckCosmo
1250 – 2500
Choosing a larger traceback-radius increases the volume of the minimal bounding ellipsoid. The exact growth of the Lagrangian ellipsoid depends on the mass and the environment of the halo. To analyze the mass growth of the proto-halo patches at different traceback radii, we compare halos from four mass bins, detailed in Table 2. Additionally, we split the $10^{11}$ $h^{-1}\mathrm{M}_{\odot}$ and $10^{13}$ $h^{-1}\mathrm{M}_{\odot}$ bins into three different environmental regimes, determined by the distance to the closest more massive halo $D_{1,1}$ (see Table 3). The results are shown in Figure 4.
3.3 Biased target selection by efficiency constraints
In order to measure these effects, we select halos with Lagrangian efficiencies below the 10th and above the 90th percentile from the $M_{\mathrm{vir}}\in[2,4]\times 10^{12}$ $h^{-1}\mathrm{M}_{\odot}$ mass bin of the 300MPC simulation at $z=0$. We then compute the distributions of a selection of external and intrinsic parameters and compare it to the distribution of all halos in that mass bin. The chosen parameters include the normalized distance to the closest more massive halo $D_{1,1}$, the relative mass increase between $1$ and $10R_{\mathrm{vir}}$ ($M_{10,1}$), the concentration parameter $c_{\mathrm{vir}}$, the spin parameter $\lambda_{\mathrm{Peebles}}$ (cf. Table 3), and the ellipticity of the minimum bounding ellipsoid $e_{\mathrm{ellipsoid}}=(a_{3}-a_{1})/(2\sum a_{i})$, where $a_{i}=\lambda_{i}^{-1/2}$ are the semi-axes of the ellipsoid related to the eigenvalues $\lambda_{1}\geq\lambda_{2}\geq\lambda_{3}$ of the shape matrix (cf. Equation 2). This definition of ellipticity is the one of Bardeen et al. (1986). The first two parameters measure the local clustering, while the following two are intrinsic properties of the halo. The ellipticity measures how spherical (and therefore well-behaved) the proto-halo patch is. Merger events and strong tidal fields are expected to distort the proto-halo patch, leading to a more elliptical and less efficient bounding ellipsoid.
4 Database architecture and web application
@@ 4.1 Architecture
Finally, the cosmicweb-music command-line interface (CLI) is a convenient tool that allows users to download initial condition configuration files directly via the command line. By using the CLI, users can avoid the cumbersome process of downloading files locally and then re-uploading them to a headless compute cluster.
For the interaction between these four components to function seamlessly, requests to and responses from the API servers must adhere to a specified protocol, which we further discuss in Appendix B. More detailed and up-to-date documentation is also available directly on the cosmICweb website [^6] https://cosmicweb.eu/documentation/api. How these requirements are best implemented on the API server side needs to be evaluated on a case-by-case basis, considering factors such as the available data, its storage, and the server infrastructure. We discuss the implementation of the API server used for the “primary” simulations in more detail at the end of this section.
4.2 Web Interface
@@
Table 3: List of currently available selectors, grouped by halo properties, galaxy properties, and further parameters including properties of available proto-halo patches. [L] and [E] mark the features that are currently supported in the locally hosted “primary” simulations and the Eagle simulations.
@@
halo $M_{200,c}$ / $M_{500,c}$ / $M_{2500,c}$ $h^{-1}\mathrm{M}_{\odot}$ Mass within $\rho\geq\Delta\rho_{c}$ contrast
@@
[…]
@@
$D_{1,1}=\|\bm{x}_{h}-\bm{x}_{n}\|/R_{h}$ ($n$: neighbor)
last major merger redshift of last major merger (with a mass ratio of 1:3 or greater)
@@
$\epsilon$ Efficiency of the Lagrangian minimum bounding ellipsoid
$\epsilon=n_{\mathrm{part}}/V_{\mathrm{ellipsoid}}$
    The primary simulations currently do not support $M_{500,b}$, $M_{2500,b}$ and the corresponding radii and concentration parameter.
  • $\diamond$
    The primary simulations use the scale factor $R_{s}$ calculated by Rockstar from $M_{\mathrm{vir}}$ and $v_{\mathrm{max}}$ (Behroozi et al., 2013a). Note that for concentrations other than $c_{\mathrm{vir}}$, there is a systematic shift in the concentrations derived from this $R_{s}^{\mathrm{vir}}$ compared to the $R_{s}$ calculated from the corresponding overdensity definition directly, see Mansfield & Kravtsov (2020).
  • $\dagger$
    The Eagle simulations currently support $M_{\mathrm{200,b}}$, $M_{\mathrm{200,c}}$, $M_{\mathrm{500,c}}$, $M_{\mathrm{2500},c}$, and $R_{\mathrm{200,c}}$.
  • $\ddagger$
    $J$ corresponds to the magnitude of the halo angular momentum and $E$ to the total energy of the halo.
Once a query is submitted, the API server returns a list of matching halos. The properties of these halos can be displayed in a table and visualized on two-dimensional scatter plost. From the results panel, users can open halos in the Halo Explorer for a more detailed overview, group and store them as collections, and directly download as initial condition files for resimulation.
@@
4.3 API implementation for the cosmICweb datasets
@@
5 Example Simulations
[…]
@@ 5 Example Simulations
Note that music allows to specify a minimum refinement level for the entire simulation, $l_{\mathrm{min}}$, which sets the number and mass of the coarsest particles in the zoom simulation, and a minimum refinement level when generating the initial conditions, $l_{\mathrm{min}}^{\mathrm{TF}}$. The particle displacements and velocities are later averaged down to $l_{\mathrm{min}}$. We present simulations varying both. This tests the dependence on the coupling of small-scale modes into the high-res regions, i.e. the accuracy of the zoom approach itself as well as its implementation in music, and the impact of a coarsely represented tidal field on the non-linear evolution during the simulation.
5.1 Milky-Way sized halo
@@
Figure 9: Visual comparison of the Au6 halo (Grand et al., 2017) between the reference simulation and three zoom simulations (cf. Figure 10). We calculate the surface density using a Delaunay tessellation field estimator (Rangel et al., 2016). The white circle indicates $R_{\mathrm{200,c}}$ determined in the reference simulation.
[…]
@@
(7,8,10) (7,10,10) (7,10,12) Ref.
$m_{p}$ [$10^{5}\mathrm{M}_{\odot}$] 364 364 5.69 2.34
$\epsilon_{\rm comoving}$ [$\mathrm{kpc}$] 3.26 3.26 0.813 0.739
$M_{200,c}$ [$10^{10}\mathrm{M}_{\odot}$] 105.3 108.1 105.2 105.4
$M_{500,c}$ [$10^{10}\mathrm{M}_{\odot}$] 83.2 83.3 83.0 82.0
$M_{2500,c}$ [$10^{10}\mathrm{M}_{\odot}$] 48.0 48.0 49.8 48.5
@@
The first example is the Au6 halo from the Auriga Project (Grand et al., 2017, 2024). This Milky Way mass halo has previously been studied to analyze formation mechanisms of disc galaxies. We perform gravity only zoom simulations at different levels of refinement of the Au6 Lagrangian patch, calculated at a traceback radius $R_{\mathrm{tb}}=4R_{\mathrm{200,c}}$ at redshift 0. The halo can be found in the publication section on cosmICweb  [^8] Auriga 6: https://cosmicweb.eu/publication/auriga-halos.
[…]
@@
Regarding global properties, we find that the halo mass $M_{\mathrm{200,c}}$ is recovered at the level of better than 1 per cent for two re-simulations, while the third, $(7,10,10)$, is somewhat more discrepant at $\sim 3$ per cent. The virial radius always agrees to better than one per cent. Properties of the inner mass distribution (concentration $c_{200,c}$ and mass $M_{2500,c}$) of the halos are slightly more discrepant but still consistent at the few per cent level. Note that the rockstar measurement of the mass and other properties of the reference run is slightly discrepant from the values reported in Grand et al. (2024) using subfind.
@@
The visual impression from Figure 9 is similar in the overall shape of the halo, also positions of some halos outside the virial radius are in excellent agreement. Interestingly, the most massive substructure is however at different locations in all of the runs — likely a consequence of it being very sensitive to any perturbations of the initial conditions. Such sensitivities are expected since orbits in virialised halos can be chaotic (see e.g. Vogelsberger et al. (2008); Sousbie & Colombi (2016); Genel et al. (2019) for some degree of evidence for the presence of chaotic mixing) implying that even small perturbations can have impact on fine-grained properties of halos, such as specifically the exact positions of substructures (see e.g. also Wang et al. (2021)).
@@
Finally, the density profiles shown in Figure 10 show agreement at the few per cent level at all radii up to the radius affected by finite force and mass resolution and out to the maximum radius of the zoom region of $4R_{200,c}$. The spherically averaged density profile is well reproduced despite the actual positions of the subhalos differing between simulations. This is possible simply because the mass associated with these substructures is small.
5.2 Group-sized halo